# TFM Dataset: A Novel Multi-task Dataset and Integrated Pipeline for Automated Tear Film Break-Up Segmentation

### v1

@@ III Tear Film Multi-task Dataset
We acquired 15 high-resolution tear film videos captured using a placido ring-based dry eye instrument. A total of 6,247 frames were extracted from these videos. Each frame was annotated for three high-level vision tasks: image classification (’Clear’, ’Closed’, ’Broken’ and ’Blur’), object detection (for placido rings and pupil region), and semantic segmentation(for tear film break-up area). Specifically, classification labels were assigned to all images. For all non-Closed frames (i.e., those displaying the placido ring), bounding boxes were annotated to localize the placido ring. Furthermore, all images labeled as ”broken” underwent pixel-level segmentation annotation for the TFBU area. The dataset composition and annotations are shown in Fig. 1. All annotations are provided using X-AnyLabeling[26].
@@ IV-B TFBU Segmentation Problem and TF-Net
Unet mobileone s0 8.6/5.4 18.9/14.4 21.8/4.6 0.652 0.731 0.973 400.3 133.6
Unet++ mobileone s0 9.9/6.7 45.4/41.0 9.8/2.4 0.610 0.684 0.969 579.3 209.8
Deeplabv3 mobileone s0 12.8/9.6 52.1/38.3 14.0/2.8 0.472 0.527 0.933 548.3 257.7
Deeplabv3+ mobileone s0 6.1/2.9 9.7/4.6 52.8/5.5 0.584 0.651 0.952 474.2 173.6
FPN mobileone s0 6.3/3.1 14.1/9.7 39.6/5.1 0.597 0.669 0.949 374.2 132.3
PAN mobileone s0 4.5/1.3 7.2/2.2 69.6/7.9 0.494 0.498 0.501 1001.3 798.1
PSPNet mobileone s0 4.4/1.2 3.4/1.3 114.7/14.9 0.654 0.735 0.973 294.5 82.3
MANet mobileone s0 36.2/33.0 24.7/20.2 17.2/4.0 0.581 0.649 0.958 631.0 243.6
LinkNet mobileone s0 5.8/2.6 8.7/4.3 49.9/6.3 0.584 0.653 0.972 486.3 190.5
UPerNet mobileone s0 14.3/11.1 58.9/54.5 11.1/2.8 0.630 0.705 0.956 417.6 144.1
Segformer mobileone s0 5.0/1.8 12.6/8.1 28.2/4.5 0.672 0.752 0.969 344.7 102.5
TF-Net (Ours) mobileone-mini0 0.72/0.65 2.7/2.4 92.2/20.0 0.698 0.779 0.975 212.9 54.1
Unet mobileone s1 9.2/9.1 19.9/19.5 18.0/3.5 0.645 0.725 0.977 413.8 129.7
Unet++ mobileone s1 12.3/12.3 77.5/77.2 6.6/1.7 0.636 0.714 0.978 432.6 141.7
Deeplabv3 mobileone s1 14.0/13.9 56.6/55.9 9.4/2.1 0.537 0.596 0.959 549.3 220.2
Deeplabv3+ mobileone s1 5.7/5.6 8.4/8.1 36.0/4.4 0.582 0.650 0.958 471.3 185.6
FPN mobileone s1 5.7/5.6 13.2/12.9 30.4/4.2 0.612 0.687 0.968 350.1 118.3
PAN mobileone s1 3.9/3.9 6.1/5.7 41.5/5.7 0.535 0.583 0.799 496.5 217.3
PSPNet mobileone s1 3.8/3.7 2.7/2.4 80.5/11.8 0.668 0.747 0.915 328.6 88.9
MANet mobileone s1 53.3/53.2 31.0/30.7 12.8/3.1 0.626 0.702 0.973 478.0 167.6
LinkNet mobileone s1 6.2/6.1 10.7/10.4 32.3/4.7 0.560 0.617 0.816 612.1 268.1
UPerNet mobileone s1 14.5/14.5 58.2/57.8 10.2/2.7 0.649 0.725 0.953 404.8 136.5
Segformer mobileone s1 4.4/4.4 11.9/11.6 22.8/3.9 0.665 0.746 0.963 275.5 79.1
TF-Net (Ours) mobileone-mini1 2.6/2.3 8.9/7.9 51.5/10.3 0.711 0.792 0.976 182.0 47.8
Unet mobileone s2 13.6/13.5 24.8/24.4 15.7/3.2 0.655 0.735 0.975 351.9 106.9
Unet++ mobileone s2 18.1/18.0 89.1/88.7 5.8/1.5 0.621 0.697 0.974 461.0 163.4
Deeplabv3 mobileone s2 22.0/21.9 88.6/87.8 6.4/1.5 0.591 0.662 0.975 387.3 130.0
Deeplabv3+ mobileone s2 9.0/8.9 12.3/11.9 27.3/3.9 0.616 0.691 0.961 396.8 129.8
FPN mobileone s2 8.2/8.2 15.9/15.5 26.0/3.8 0.625 0.700 0.950 329.2 103.7
PAN mobileone s2 6.4/6.3 9.2/8.8 30.4/5.2 0.525 0.549 0.562 471.1 228.8
PSPNet mobileone s2 6.2/6.1 3.8/3.6 63.4/10.1 0.666 0.748 0.966 292.7 79.1
MANet mobileone s2 126.1/126.1 52.6/52.2 8.5/2.3 0.636 0.712 0.967 441.6 151.9
LinkNet mobileone s2 12.1/12.0 17.4/17.0 24.6/4.2 0.624 0.701 0.973 428.5 149.4
UPerNet mobileone s2 19.4/19.3 61.2/60.9 9.5/2.5 0.676 0.753 0.954 337.1 104.0
Segformer mobileone s2 7.0/6.9 14.6/14.2 20.6/3.6 0.668 0.749 0.967 304.1 81.3
TF-Net (Ours) mobileone-mini2 5.7/5.2 19.8/17.6 24.5/6.1 0.720 0.800 0.973 180.8 47.0
Unet mobileone s3 16.3/16.2 28.9/28.5 13.4/2.9 0.656 0.737 0.977 397.0 113.4
Unet++ mobileone s3 22.9/22.8 113.6/113.2 4.9/1.3 0.631 0.702 0.834 549.9 190.6
Deeplabv3 mobileone s3 24.3/24.2 98.1/97.2 5.7/1.4 0.611 0.686 0.974 366.4 113.5
Deeplabv3+ mobileone s3 11.3/11.2 15.7/15.2 22.1/3.5 0.647 0.726 0.961 292.5 84.9
FPN mobileone s3 10.6/10.5 19.4/18.9 21.4/3.6 0.636 0.714 0.962 305.5 96.0
PAN mobileone s3 8.7/8.6 12.9/12.4 23.7/4.5 0.521 0.570 0.808 613.0 314.5
PSPNet mobileone s3 8.6/8.5 5.4/5.1 50.6/8.4 0.706 0.786 0.952 234.7 60.4
MANet mobileone s3 129.5/129.4 57.7/57.3 7.7/1.0 0.661 0.741 0.978 367.4 108.6
LinkNet mobileone s3 14.7/14.6 22.8/22.3 19.4/3.6 0.636 0.714 0.973 378.6 123.6
UPerNet mobileone s3 21.8/21.7 64.7/64.2 8.8/2.4 0.657 0.734 0.952 360.0 123.3
Segformer mobileone s3 9.3/9.2 18.0/17.6 17.5/3.1 0.684 0.765 0.970 242.7 70.3
TF-Net (Ours) mobileone-mini3 10.2/9.2 35.1/31.3 17.8/4.2 0.721 0.801 0.972 187.5 51.9
Unet mobileone s4 21.5/21.4 36.2/35.6 8.7/2.4 0.664 0.744 0.975 361.1 101.7
Unet++ mobileone s4 33.4/33.3 171.4/170.8 3.0/0.9 0.673 0.753 0.977 331.6 92.7
Deeplabv3 mobileone s4 29.0/28.9 111.4/110.2 3.0/1.1 0.583 0.653 0.967 420.5 143.2
Deeplabv3+ mobileone s4 16.1/16.0 21.8/21.1 11.0/2.8 0.629 0.705 0.963 320.2 108.1
FPN mobileone s4 15.4/15.3 25.6/25.0 11.5/3.2 0.644 0.724 0.964 285.7 83.6
PAN mobileone s4 13.6/13.5 19.4/18.8 11.2/3.3 0.527 0.566 0.785 710.7 395.1
PSPNet mobileone s4 13.6/13.5 9.6/9.1 34.3/6.4 0.664 0.746 0.965 310.7 83.4
MANet mobileone s4 135.7/135.5 66.7/66.1 5.8/1.8 0.629 0.704 0.975 423.7 153.0
LinkNet mobileone s4 20.0/19.9 33.1/32.5 10.5/2.9 0.647 0.727 0.979 364.0 120.7
UPerNet mobileone s4 26.6/26.5 70.9/70.3 6.5/2.1 0.642 0.720 0.942 368.3 130.2
Segformer mobileone s4 14.2/14.1 24.3/23.7 10.2/2.3 0.670 0.750 0.973 286.4 83.3
TF-Net (Ours) mobileone-mini4 15.9/14.4 54.6/48.8 10.3/3.0 0.737 0.814 0.965 175.6 44.8
TABLE II: Binary segmentation performance comparison of various models on the TF-Crop test set. Our TF-Net utilizes MobileOne-mini backbone - a further streamlined version of the original MobileOne[25] architecture where stage-wise block repetitions are reduced from [2, 8, 10, 1] to [2, 3, 4, 3]. Additionally, we strategically constrain the channel dimensions across network layers to further minimize parameter count and computational complexity. These architectural refinements collectively enable substantially accelerated inference speeds on mobile platforms while maintaining competitive segmentation accuracy.
@@ V-B Comparative Experiments
As presented in Table II, alongside standard medical segmentation metrics, we specifically compared model parameters, FLOPs, and FPS(frames per second) on both a GPU (RTX 3050 Ti) and a CPU (Intel Core i5). This facilitates a holistic assessment of both segmentation accuracy and practical deployment potential. The results clearly demonstrate that our proposed TF-Net not only surpasses all other models across various segmentation metrics but also achieves significantly faster inference speeds, attaining a processing rate of 92.2(on GPU)/20.0(on CPU) frames per second. This level of performance adequately fulfills the requirements for real-time clinical diagnosis.
@@ V-D System Robustness and TF-Collab Evaluation
The pipeline successfully processed input videos in real-time, achieving an average throughput of 92.2(on GPU, 3050ti)/20.0(on CPU, core i5) frames per second. Its end-to-end segmentation performance for frames identified as Broken yielded a mDSC of 0.779, which is only marginally lower than TF-Net’s standalone performance on the curated Broken-only test set. This minor performance drop can be attributed to occasional inaccuracies in the placido rings detection, leading to sub-optimal cropping, rather than failures of the segmentation core itself.

### v2

@@ III Tear Film Multi-task Dataset
We acquired multiple high-resolution tear film videos captured using a Placido ring-based dry eye instrument. A total of 6,247 valid frames were extracted from these videos. Each frame was annotated for three high-level vision tasks: image classification (’Clear’, ’Closed’, ’Broken’, and ’Blur’), object detection (for Placido rings and pupil region), and semantic segmentation (for tear film break-up area). Specifically, classification labels were assigned to all images. For all non-Closed frames (i.e., those displaying the Placido ring), bounding boxes were annotated to localize the Placido ring. Furthermore, all images labeled as ”Broken” underwent pixel-level segmentation annotation for the TFBU area. The dataset composition and annotations are shown in Fig. 1. All annotations are provided using X-AnyLabeling[26].
@@ IV-B TFBU Segmentation Problem and TF-Net
Unet mobileone s0 8.6/5.4 18.9/14.4 76.1/4.6 0.652 0.731 0.973 400.3 133.6
Unet++ mobileone s0 9.9/6.7 45.4/41.0 32.0/2.4 0.610 0.684 0.969 579.3 209.8
Deeplabv3 mobileone s0 12.8/9.6 52.1/38.3 54.6/2.8 0.472 0.527 0.933 548.3 257.7
Deeplabv3+ mobileone s0 6.1/2.9 9.7/4.6 168.3/5.5 0.584 0.651 0.952 474.2 173.6
FPN mobileone s0 6.3/3.1 14.1/9.7 124.3/5.1 0.597 0.669 0.949 374.2 132.3
PAN mobileone s0 4.5/1.3 7.2/2.2 207.2/7.9 0.494 0.498 0.501 1001.3 798.1
PSPNet mobileone s0 4.4/1.2 3.4/1.3 319.8/14.9 0.654 0.735 0.973 294.5 82.3
MANet mobileone s0 36.2/33.0 24.7/20.2 59.7/4.0 0.581 0.649 0.958 631.0 243.6
LinkNet mobileone s0 5.8/2.6 8.7/4.3 131.6/6.3 0.584 0.653 0.972 486.3 190.5
UPerNet mobileone s0 14.3/11.1 58.9/54.5 44.3/2.8 0.630 0.705 0.956 417.6 144.1
Segformer mobileone s0 5.0/1.8 12.6/8.1 57.6/4.5 0.672 0.752 0.969 344.7 102.5
TF-Net (Ours) mobileone-mini0 0.72/0.65 2.7/2.4 302.2/33.8 0.698 0.779 0.975 212.9 54.1
Unet mobileone s1 9.2/9.1 19.9/19.5 61.7/3.5 0.645 0.725 0.977 413.8 129.7
Unet++ mobileone s1 12.3/12.3 77.5/77.2 21.9/1.7 0.636 0.714 0.978 432.6 141.7
Deeplabv3 mobileone s1 14.0/13.9 56.6/55.9 36.6/2.1 0.537 0.596 0.959 549.3 220.2
Deeplabv3+ mobileone s1 5.7/5.6 8.4/8.1 115.1/4.4 0.582 0.650 0.958 471.3 185.6
FPN mobileone s1 5.7/5.6 13.2/12.9 96.0/4.2 0.612 0.687 0.968 350.1 118.3
PAN mobileone s1 3.9/3.9 6.1/5.7 130.7/5.7 0.535 0.583 0.799 496.5 217.3
PSPNet mobileone s1 3.8/3.7 2.7/2.4 222.2/11.8 0.668 0.747 0.915 328.6 88.9
MANet mobileone s1 53.3/53.2 31.0/30.7 43.2/3.1 0.626 0.702 0.973 478.0 167.6
LinkNet mobileone s1 6.2/6.1 10.7/10.4 87.4/4.7 0.560 0.617 0.816 612.1 268.1
UPerNet mobileone s1 14.5/14.5 58.2/57.8 40.0/2.7 0.649 0.725 0.953 404.8 136.5
Segformer mobileone s1 4.4/4.4 11.9/11.6 50.8/3.9 0.665 0.746 0.963 275.5 79.1
TF-Net (Ours) mobileone-mini1 2.6/2.3 8.9/7.9 166.5/17.5 0.711 0.792 0.976 182.0 47.8
Unet mobileone s2 13.6/13.5 24.8/24.4 53.4/3.2 0.655 0.735 0.975 351.9 106.9
Unet++ mobileone s2 18.1/18.0 89.1/88.7 19.6/1.5 0.621 0.697 0.974 461.0 163.4
Deeplabv3 mobileone s2 22.0/21.9 88.6/87.8 25.6/1.5 0.591 0.662 0.975 387.3 130.0
Deeplabv3+ mobileone s2 9.0/8.9 12.3/11.9 87.6/3.9 0.616 0.691 0.961 396.8 129.8
FPN mobileone s2 8.2/8.2 15.9/15.5 80.7/3.8 0.625 0.700 0.950 329.2 103.7
PAN mobileone s2 6.4/6.3 9.2/8.8 95.3/5.2 0.525 0.549 0.562 471.1 228.8
PSPNet mobileone s2 6.2/6.1 3.8/3.6 181.2/10.1 0.666 0.748 0.966 292.7 79.1
MANet mobileone s2 126.1/126.1 52.6/52.2 29.1/2.3 0.636 0.712 0.967 441.6 151.9
LinkNet mobileone s2 12.1/12.0 17.4/17.0 68.6/4.2 0.624 0.701 0.973 428.5 149.4
UPerNet mobileone s2 19.4/19.3 61.2/60.9 36.4/2.5 0.676 0.753 0.954 337.1 104.0
Segformer mobileone s2 7.0/6.9 14.6/14.2 46.2/3.6 0.668 0.749 0.967 304.1 81.3
TF-Net (Ours) mobileone-mini2 5.7/5.2 19.8/17.6 83.2/9.4 0.720 0.800 0.973 180.8 47.0
Unet mobileone s3 16.3/16.2 28.9/28.5 46.7/2.9 0.656 0.737 0.977 397.0 113.4
Unet++ mobileone s3 22.9/22.8 113.6/113.2 16.6/1.3 0.631 0.702 0.834 549.9 190.6
Deeplabv3 mobileone s3 24.3/24.2 98.1/97.2 22.6/1.4 0.611 0.686 0.974 366.4 113.5
Deeplabv3+ mobileone s3 11.3/11.2 15.7/15.2 71.7/3.5 0.647 0.726 0.961 292.5 84.9
FPN mobileone s3 10.6/10.5 19.4/18.9 68.0/3.6 0.636 0.714 0.962 305.5 96.0
PAN mobileone s3 8.7/8.6 12.9/12.4 77.5/4.5 0.521 0.570 0.808 613.0 314.5
PSPNet mobileone s3 8.6/8.5 5.4/5.1 147.0/8.4 0.706 0.786 0.952 234.7 60.4
MANet mobileone s3 129.5/129.4 57.7/57.3 26.5/1.0 0.661 0.741 0.978 367.4 108.6
LinkNet mobileone s3 14.7/14.6 22.8/22.3 56.9/3.6 0.636 0.714 0.973 378.6 123.6
UPerNet mobileone s3 21.8/21.7 64.7/64.2 33.5/2.4 0.657 0.734 0.952 360.0 123.3
Segformer mobileone s3 9.3/9.2 18.0/17.6 41.7/3.1 0.684 0.765 0.970 242.7 70.3
TF-Net (Ours) mobileone-mini3 10.2/9.2 35.1/31.3 64.3/4.2 0.721 0.801 0.972 187.5 51.9
Unet mobileone s4 21.5/21.4 36.2/35.6 32.5/2.4 0.664 0.744 0.975 361.1 101.7
Unet++ mobileone s4 33.4/33.3 171.4/170.8 11.4/0.9 0.673 0.753 0.977 331.6 92.7
Deeplabv3 mobileone s4 29.0/28.9 111.4/110.2 12.9/1.1 0.583 0.653 0.967 420.5 143.2
Deeplabv3+ mobileone s4 16.1/16.0 21.8/21.1 41.9/2.8 0.629 0.705 0.963 320.2 108.1
FPN mobileone s4 15.4/15.3 25.6/25.0 42.4/3.2 0.644 0.724 0.964 285.7 83.6
PAN mobileone s4 13.6/13.5 19.4/18.8 43.0/3.3 0.527 0.566 0.785 710.7 395.1
PSPNet mobileone s4 13.6/13.5 9.6/9.1 106.6/6.4 0.664 0.746 0.965 310.7 83.4
MANet mobileone s4 135.7/135.5 66.7/66.1 21.0/1.8 0.629 0.704 0.975 423.7 153.0
LinkNet mobileone s4 20.0/19.9 33.1/32.5 36.0/2.9 0.647 0.727 0.979 364.0 120.7
UPerNet mobileone s4 26.6/26.5 70.9/70.3 25.9/2.1 0.642 0.720 0.942 368.3 130.2
Segformer mobileone s4 14.2/14.1 24.3/23.7 30.4/2.3 0.670 0.750 0.973 286.4 83.3
TF-Net (Ours) mobileone-mini4 15.9/14.4 54.6/48.8 38.3/3.0 0.737 0.814 0.965 175.6 44.8
TABLE II: TFBU segmentation performance comparison of various models on the TF-Crop test set. Our TF-Net utilizes MobileOne-mini backbone - a further streamlined version of the original MobileOne[25] architecture where stage-wise block repetitions are reduced from [2, 8, 10, 1] to [2, 3, 4, 3]. Additionally, we strategically constrain the channel dimensions across network layers to further minimize parameter count and computational complexity. These architectural refinements collectively enable substantially accelerated inference speeds on mobile platforms while maintaining competitive segmentation accuracy.
@@ V-B Comparative Experiments
As presented in Table II, alongside standard medical segmentation metrics, we specifically compared model parameters, FLOPs, and FPS(frames per second) on both a GPU (RTX 3050 Ti) and a CPU (Intel Core i5). This facilitates a holistic assessment of both segmentation accuracy and practical deployment potential. The results clearly demonstrate that our proposed TF-Net not only surpasses all other models across various segmentation metrics but also achieves significantly faster inference speeds, attaining a processing rate of 318.2(on GPU)/33.8(on CPU) frames per second. This level of performance adequately fulfills the requirements for real-time clinical diagnosis.
@@ V-D System Robustness and TF-Collab Evaluation
The pipeline successfully processed input videos in real-time, achieving an average throughput of 318.2(on GPU, 3050ti)/33.8(on CPU, core i5) frames per second. Its end-to-end segmentation performance for frames identified as Broken yielded a mDSC of 0.779, which is only marginally lower than TF-Net’s standalone performance on the curated Broken-only test set. This minor performance drop can be attributed to occasional inaccuracies in the placido rings detection, leading to sub-optimal cropping, rather than failures of the segmentation core itself.